OpenStudy (anonymous):

WHAT IS THE SUM OF ALL EVEN INTERGERS BETWEEN 99 AND 301

OpenStudy (ajprincess):

Even numbers between 99 and 301 are 100,102................,300. This an arithmetic progression with first term 100, common difference 2 and the last term 300. Using \(a_n=a+(n-1)d\) \(300=100+(n-1)2\)n can be found. To find the sum use the formula\[S_n=\frac{ n }{2}(a+l)\] a-first term l-last term.